• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Mohamed

    Log entry time 19:39:19 on May 08, 2010

    Entry number 321773

    keyword=Compton cavity status

    - Heating the cavity has been terminated at 2:00 PM. The pressure decreased to mid of 10**-6 Torr. Then, the cavity has been Helium leak-checked. Following, the cavity is pumped with the ion pumps. - The resonance mode has been checked and the cavity mirrors have been re-aligned to improve the mode. While the cavity was connected to the turbo pump, the reflection signal showed ~70%. - The gate valve VBV1P02 work automatically after the pressure improved to low 10**-7 Torr. - The system is under ion pumps and pressure showed 1x10**-7 after ~30 minutes, expecting be in the 10**-8 range after few hours.



    A copy of this log entry has been emailed to: nanda@jlab.org